OpenStudy (anonymous):

WILL FAN AND MEDAL!! what is the slope of the line passing through (2,7) and (-1, 3)

OpenStudy (anonymous):

y2-y1 3-7=-4 -1-2=-3 slope = -4/3

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

your welcome! :-)